OpenStudy (anonymous):

use incepet from method to graph fom. here the problem is 30x+y=-30

OpenStudy (anonymous):

what do you think you'd do?

OpenStudy (anonymous):

you get y bye its self

OpenStudy (anonymous):

and how would you think to do that?

OpenStudy (anonymous):

we only have 30 in the number

OpenStudy (anonymous):

30X+y=-30 We have to think of something that equals -30..?

OpenStudy (anonymous):

What does the X represent

OpenStudy (anonymous):

5

OpenStudy (anonymous):

how did you get 5?

OpenStudy (anonymous):

t goses in to it

OpenStudy (anonymous):

alright. now what do you think you'd do next?

OpenStudy (anonymous):

so u dive the 30x by 5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but first did you find out what X is? example of how to find the X:x - 4 = 10

OpenStudy (anonymous):

that's not torward this problem just showing an example to find your X

OpenStudy (anonymous):

now i have 6x+y =-6

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

6 minus what equals NEGATIVE six?

OpenStudy (anonymous):

what's six plus 6?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Now subtract 6 - 12. what do you get?

OpenStudy (anonymous):

no its 12

OpenStudy (anonymous):

6+6= 12

OpenStudy (anonymous):

now whats 6-12?

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so what do you think is our Y?

OpenStudy (anonymous):

? -1

OpenStudy (anonymous):

no........

OpenStudy (anonymous):

can u help -6

OpenStudy (anonymous):

do you understand what I am trying to explain to you?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

first we started of doing 6+6=12))) then we did 6-12 what did we get..? -6. So our missing number that goes to Y is 12. we used that to solve the problem 6x-y(12=-6

OpenStudy (anonymous):

because i went 30 dived bye -6